OpenStudy (anonymous):

Of 50 employees of a store located in downtown Boston, 18 people take the subway to work, 12 take the bus, and 7 take both the subway and the bus. How may employees (a) take the subway or the bus to work? (b) take only the bus to work? (c) take either the bus or the subway to work? (d) get to work by some other means?

OpenStudy (anonymous):

n(P) = 50 (total no. of employees) n(S) = 18 (no. of people taking subway) n(B) = 12 (no. of people taking the bus) N(S x B) = 7 (taking both subway and bus) a. take subway or bus = N(S) + N(B) - N(S x B) = 18 + 12 - 7 = 23 b. take only bus = N(B) - n(S x B) = 12 - 7 = 5 c. take either bus or subway = same as eqn a d. get to work by other means = n(P) - eqn (a) = 50 - 23 = 27